Cisplatin or Platinol has a chemical PtCl2(NH3)2 that is used in the treatment of certain cancers. Calculate (a) the amount (mol) of compound in 285.3 g of cisplatin


1
Expert's answer
2022-03-17T14:57:00-0400

Molar mass of cisplatin "[Pt(NH_3)_2Cl_2]=301.1 \frac{g}{mol}"


mass of cisplatin is given as 285.3 g


Mole of cisplatin "= \frac{mass}{molarmass}= \frac{285.3}{301.1}=0.9475" mole